OpenStudy (anonymous):

Slope=5, passing through (-4,8) answer in Slope-I form s/b: y=5x+28. my answer y=5x+-12 ???

OpenStudy (anonymous):

using y - y0=m(x-x0) we get: x0=-4 y0=8 y-8 = 5(x-(-4)) y=5x+28